Shift report 20000728 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Fri Jul 28 2000 - 15:52:37 EDT

  • Next message: http account: "Shift report 20000728 16:00-24:00"

    ------------------------------------------------------------
    Date: Fri Jul 28 15:52:37 2000, Shift: 08:00-16:00
    Supervisor: kh, with: +oh
    Log book#3 pages 102 - 105
    
    Shift Summary:
    ------------------------------------------------------------
    1. No beam, playing with no. of bunches in blue ring
    2. One pedestal run, 1967. Excercised field in D5 and set
       D1 to 520 A. First counting house meeting conducted.
    3. Some trips in MULT Si (13.0,13.8,9,10) reset.
    4. Run plan for weekend
       a) Run 10k triggers with following conditions:
          i)   D1 @ 520A (1/6 field)
          ii)  D5 off
          iii) T1 gating grid set to 180+-80 (already set)
          iv)  MTPC1 gating grid set to 200+-100 (already set)
          v)   ZDC with high threshold set at 50mV (already set)
    
       b) Run 10k triggers with following conditions:
          i)   D5 @ 750A (1/4 field)
          ii)  other conditions same as previous run
       c) Run 10k triggers with following conditions:
          i)   T1 gating grid off
          ii)  MTPC1 gating grid set to 250 +- 80
          iii) ZDC threshold turned to 100mV
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Fri Jul 28 2000 - 15:57:39 EDT